OpenStudy (anonymous):

Suppose that we know that f(x) is a continuous and differentiable function on [2,14]. Also, suppose we know that f(2)=9 and f'(x) is greater than or equal to -12. What is the smallest possible value of f(14)?

OpenStudy (amistre64):

would this be a straight line from 2,14 running at a slope of 12?

OpenStudy (anonymous):

(I'm supposed to use the mean value theorem, or at least that's what it says on the worksheet.)

OpenStudy (anonymous):

(I'm supposed to use the mean value theorem, or at least that's what it says on the worksheet.)

OpenStudy (anonymous):

@amistre64 - I have no idea! I'm complete lost on this one. Don't even know where to start.

OpenStudy (amistre64):

f'(2) =>12....6x? maybe

OpenStudy (anonymous):

Well, if f'(x) = -12 throughout the interval then it would be a straight line from (2,9) with a slope of -12. So that would be the minimum possible value given what we know.

OpenStudy (anonymous):

Sorry, @polpak - how do we know that?

OpenStudy (amistre64):

ahh,....(-)12 ... that was hiding in the corner

OpenStudy (anonymous):

We are given that f'(x) is greater than or equal to -12.

OpenStudy (anonymous):

So if it maintained the smallest possible derivative over the whole interval we would have a line with slope of -12.

myininaya (myininaya):

i got the smallest possible value is -135 for f(14)

OpenStudy (amistre64):

thats what I thought :)

OpenStudy (amistre64):

but with a 12 lol

OpenStudy (anonymous):

9-12*12 = 9-144 = -135

myininaya (myininaya):

yay thats what i got except differently

myininaya (myininaya):

-12<=[f(2)-f(14)]/[2-14]

OpenStudy (anonymous):

Does that make sense Chap?

OpenStudy (anonymous):

How do you figure what f(14) is?

myininaya (myininaya):

chap you can also solve the above for f(14) resulting in f(14)>=(-135)

myininaya (myininaya):

I used the mean value thm just like you said

OpenStudy (anonymous):

Right. But I have -12<= 9 - f(14)

OpenStudy (anonymous):

Oops wait

OpenStudy (anonymous):

I have -12 <= 9 - f(14) ----------- -12

myininaya (myininaya):

multipliy -12 on both sides don't forget to flip the inequality when you multipliy or divide by a negative

OpenStudy (anonymous):

Sorry )= I don't understand how to do that.

OpenStudy (anonymous):

\[-12 \le \frac{9- f(14)}{-12} \implies 144 \ge 9-f(14)\]

myininaya (myininaya):

how do you solve x/5=4

OpenStudy (anonymous):

\[-4a < b \implies a \gt \frac{b}{-4}\]

OpenStudy (anonymous):

Okay. But how does 135 >= - f(14) solve for f(14)?

OpenStudy (anonymous):

Multiply by -1

myininaya (myininaya):

now multipliy both sides by negative 1 don't forget to flip

OpenStudy (anonymous):

135 <= f(14)

myininaya (myininaya):

right!

myininaya (myininaya):

wait where's the negative?

OpenStudy (anonymous):

That's the answer?

OpenStudy (anonymous):

-135 <= f(14)

myininaya (myininaya):

ok so we have f(14)>=-135 this means the smallest that f(14) can be is -135

OpenStudy (anonymous):

OH! I thought f(14) would be positive. Thank you so much!

OpenStudy (anonymous):

It might be positive!

OpenStudy (anonymous):

We only know that the smallest it can be is -135. It could be very very large.

myininaya (myininaya):

it could be 1111111111111111111111

OpenStudy (anonymous):

over 9000!

myininaya (myininaya):

5 trillion billion if that means anything